Max spends three times as long on his math homework than he does on his science homework. if he spent a total of 64 minutes on math and science homework last night, how long did he spend on math homework


1
Expert's answer
2021-02-02T04:37:35-0500

Let m=math and s=science


m=3s

m+s=64


So using substitution,

3s+s=64

4s=64

s=16


So m=64-16=48 minutes